OpenStudy (anonymous):

a four person committee is to be elected from an organization's membership of 11 people. How many different committees are possible?

OpenStudy (aric200):

= 330

OpenStudy (anonymous):

the number of different commitees available = C (11,4) = 11 !/(4!*7!)

OpenStudy (aric200):

11 C 4 = (11!)/(4!(11-4)!) = (11!)/(4!7!) = (11*10*9*8*7!)/(4!7!) = (11*10*9*8)/(4*3*2*1) = 7920/24 = 330

OpenStudy (aric200):

Thanks for the medal :)